# Compact quantum algorithms for time-dependent differential equations

### v2

@@ IV.1 TMCQC1 - Explicit Expansion Circuit
$\implies\sum_{|v|=\tau}{\tau\choose v}U^{v}\equiv\sum_{k=0}^{N_{T}-1}\tilde{U}_{k},$ (15)
where $v=(v_{1},\cdots,v_{K})$ and $U^{v}=U^{v_{1}}_{0}U^{v_{2}}_{1}\cdots U^{v_{K}}_{K-1}$.
@@ IV.2 TMCQC2 - Explicit Serial Circuit
The previous approach can be improved by trading in a few extra qubits, while still keeping the total qubit complexity logarithmic in the problem size. An alternative way to implement the same operation as earlier is to concatenate the LCU decomposition of $A_{E}$ serially, $\tau$ times. However, to keep track of the time steps, we would need an additional register, which we shall refer to as clock register, comprising $n_{c}=\log_{2}(\tau)$ qubits. The quantum circuit for this is shown in figure 2(a) (for $K=4)$. A step-by-step description of this circuit operation is outlined in Section LABEL:subsec:_Explicit_TMCQC of the Appendix.
@@ IV.3 TMCQC3 - Implicit Expansion Circuit
Gate complexity - To compute the gate complexity, first consider a $p$th order term in the truncated series. If this term is written in terms of the unitaries decomposition, it will produce ($p+K-1\choose K-1$ terms, which is $\mathcal{O}(p^{K-1})$. Recall that each of these newly produced terms is in fact a product of $p$th order products (at most) of unitaries, thus making the depth $\mathcal{O}(p^{K})$. Next, we sum the series up to $P_{min}$ terms, giving a total of (for $K=2$) $\mathcal{O}(P^{3}_{min})$. Finally, since we need to perform time marching, the series obtained above is finally raised to the exponent $\tau$, expanding which gives a total depth of $N_{T}=$(${\tau+(P^{\text{min}})^{3}-1\choose(P^{\text{min}})^{3}-1}=\mathcal{O}(\tau^{P^{3}_{min}})$. Therefore the overall gate complexity is $\mathcal{O}(G_{U}\tau^{P^{3}_{min}})$. It is clear that the algorithm is exponential in $\tau$ and also clearly worse (in terms of $\tau$) than the corresponding classical complexity given by $\mathcal{O}(\log(N_{g}s\tau\kappa\log(1/\varepsilon)))$. Again, although the algorithm is still logarithmic in $N_{g}$, it offers the poorest complexity scaling compared to all other TMCQCs presented here.
@@ IV.4 TMCQC4 - Implicit Serial Circuit
Gate complexity - The truncated series has a total of $\mathcal{O}(P^{3}_{min})$ terms, so the total depth can be estimated to be $\mathcal{O}(P^{3}_{min}\tau)$. Thus it can be seen readily that the total gate complexity is $\mathcal{O}\Big{(}s\epsilon,\log(N_{g}\tau),\textrm{polylog}(\epsilon/\varepsilon_{U},1/\varepsilon_{N}),\log^{-3}((\kappa-1)^{-1})\Big{)}$. Comparing this with the classical complexity of $(\log(N_{g}s\tau\kappa\log(1/\varepsilon)))$, one finds that the algorithm (without any parallelization) has a comparable scaling in $\tau$, while it is exponentially better than classical in $N_{g}$ and $\kappa$.
@@ IV.6 Query complexity and success probability
The above value is for a single linear combination of unitaries block. However in essence, each TMCQC provides a different unitary block-encoding of the corresponding $A$ to perform time marching. The $p_{\textrm{succ}}$ of the solution at $t=\tau\Delta t$ thus varies for each TMCQC. The specific dependence arises from the difference in the number of times the unitaries oracle is queried and the corresponding depth, which we shall represent from now on as $G_{\textrm{L}}$. Comparing all the TMCQCs we can easily identify that in TMCQC2 and TMCQC4, the decay in $p_{\textrm{succ}}$ is particularly the most severe since the unitaries are applied in series. Therefore for $\tau$ such applications, the RHS of eq. (18) will now be raised to the power $\tau$. For generality, let us represent this exponent with $G_{\textrm{L}}$. A single application of the two-unitary oracle involves preparing a superposition state by Hadamard gates acting on ancilla qubits. This implies that $\beta=2$. Since the $p_{\textrm{succ}}$ for every time step is independent, the probability thus tends to decay exponentially [berry2022quantum, fang2023time], compounding for each of the $\tau$ steps. Apart from this, the overall success probability also depends on ratio of the norms of the initial and final time solution states. Thus, we can rewrite $p_{\textrm{succ}}$ as
@@
In summary, it is ideal to ensure that $\eta\lesssim 1$, otherwise it would lead to a large pre-factor $\mathcal{O}(2^{\log(\tau)/\log(\kappa-1)})$ at best, or $\mathcal{O}(2^{\tau^{2}})$ at worst. While ensuring the query complexity to be controllable, the following three situations need special attention.
@@ V End-to-end algorithm and near-term strategies
(A) Shot parallelization - Generally the shot count available on real quantum hardware (here, IBMQ) is limited to about $\sim 2^{15}\textrm{~{}to~{}}2^{19}$ shots. If we consider a simulation requiring a large shot-count of about (say) $N_{s}=2^{20}$ shots or higher, such an execution would not be possible with current hardware. A simple workaround to achieve this would be actually to execute multiple quantum circuits in parallel. For the current example, naively, we can perform two parallel circuit executions with $2^{19}$ shots each. When possible, we could go even further by executing several low shot-count circuits in parallel, with which we can increase the total shot count while improving the sampling accuracy of the quantum state. To fit our simulations using only the available shot count, we could instead use the Richardson extrapolation to lower the required $N_{s}$. This is discussed below in this section.
(B) LCU parallelization - Let us consider a single time step simulation using either two or four unitary block encoding.
@@
    Naive Parallel unitaries - We execute each unitary $U_{i}$ (without any controls) in parallel on 2 (or 4) separate circuits, starting with the same initial state. This requires $3n_{g}$ qubits in total. When the resulting parallel (partial) solution states have purely real amplitudes, we can perform a straightforward measurement in the computational basis of each partial solution. Following this, these partial solutions can be combined classically to output the final solution. However, if any of the partial solutions have complex amplitudes, a direct extraction of solutions requires expensive tomography to reconstruct the state. Two possibilities exist. In the first case, we combine those parallel unitaries as a single circuit such that the grouping produces real amplitudes for the partial states. From here, we can proceed with measurement as earlier. This new grouping would, of course, increase the depth of the circuit, which is now at some intermediate level of parallelization. Secondly, if we are interested only in computing the expectation value of a certain observable, instead of the full state, it can be done while still retaining the full parallelization. The expectation value of an operator $U_{i}$ can be computed accurately by using a Hadamard Test circuit. A direct Hadamard test outputs $Re(\langle\psi|U_{i}|\psi\rangle)$. For the imaginary part $Im(\langle\psi|U_{i}|\psi\rangle)$, a simple modification is done by adding the $S^{{\dagger}}$ gate as shown in figure 2(b).
@@
    Fanout Parallel unitaries - A more robust parallelization is possible by invoking the idea of fanout quantum circuits [nielsen2010quantum, cleve2000fast, moore2001parallel], though is is hard to realize using near-term devices. Here, the circuit can be parallelized by a single entangled quantum circuit (not separate circuits as earlier) at the cost of extra ancilla qubits [hoyer2005quantum, boyd2023low, zhang2024parallel]. In these types of strategies, the state of a single ancilla register prepared with the target state is then “basis-type copied” (fanout) to other additional ancilla registers by applying controlled CNOT gates as $|\psi\rangle|0\rangle\cdots|0\rangle\rightarrow|\psi\rangle|\psi\rangle\cdots|\psi\rangle$. Now the unitary operators are applied in parallel on copies of the target state $|\psi\rangle$. Especially when the Hamiltonian that is being simulated has certain properties [zhang2024parallel], or when the set of unitaries can be partitioned into Pauli operators, such fanout schemes can be used, with the aid of Clifford circuits [boyd2023low], to parallelize the overall algorithm.
@@
where $\epsilon_{1}>\epsilon_{2}$ and $\gamma=\epsilon_{1}/\epsilon_{2}$ is the order of extrapolation, where the error of extrapolation scales as $\mathcal{O}(\epsilon^{4})$. $\Gamma$ here represents, collectively, any arbitrary set of parameters on which the operator could depend. Higher powers of $\gamma$ leads to higher order and more accurate extrapolations. As we shall demonstrate later through simulations, even for $\epsilon_{1}and\epsilon_{2}$ close to unity, the solution can be computed accurately through extrapolation. This method could also serve as an aid to possible amplitude amplification procedures [berry2014exponential]. The extrapolation procedure can be viewed alternatively as follows. Given a fixed number of shots $N_{s}$, we can extrapolate the solutions that were obtained with $\epsilon_{1,2}$ and $N_{s}$ shots as

### v3

@@ IV.1 TMCQC1 - Explicit Expansion Circuit
$\implies\sum_{\begin{subarray}{c}v_{1}+v_{2}+\cdots+v_{k}=\tau\\ v_{i}\geq 0\forall i\end{subarray}}{\tau\choose v_{1},\cdots,v_{K}}U^{v_{1},\cdots,v_{K}}\equiv\sum_{k=0}^{N_{T}-1}\tilde{U}_{k},$ (15)
where (${\tau\choose v}=\tau!/(v_{1}!,v_{2}!\cdots,v_{K}!)$.
@@ IV.2 TMCQC2 - Explicit Serial Circuit
The previous approach can be improved by trading in a few extra qubits, while still keeping the total qubit complexity logarithmic in the problem size. An alternative way to implement the same operation as earlier is to concatenate the LCU decomposition of $A_{E}$ serially, $\tau$ times. However, to keep track of the time steps, we would need an additional register, which we shall refer to as clock register [berry2024quantum, fang2023time], comprising $n_{c}=\log_{2}(\tau)$ qubits. The quantum circuit for this is shown in figure 2(a) (for $K=4)$. A step-by-step description of this circuit operation is outlined in Section LABEL:subsec:_Explicit_TMCQC of the Appendix.
@@ IV.3 TMCQC3 - Implicit Expansion Circuit
Gate complexity - To compute the gate complexity, first consider a $p$th order term in the truncated series. If this term is written in terms of the unitaries decomposition, it will produce ($p+K-1\choose K-1$ terms, which is $\mathcal{O}(p^{K-1})$. Recall that each of these newly produced terms is in fact a product of $p$th order products (at most) of unitaries, thus making the depth $\mathcal{O}(p^{K})$. Next, we sum the series up to $P_{min}$ terms, giving a total of (for $K=2$) $\mathcal{O}(P^{3}_{min})$. Finally, since we need to perform time marching, the series obtained above is finally raised to the exponent $\tau$, expanding which gives a total depth of $N_{T}=$(${\tau+(P^{\text{min}})^{3}-1\choose(P^{\text{min}})^{3}-1}=\mathcal{O}(\tau^{P^{3}_{min}})$. Therefore the overall gate complexity is $\mathcal{O}(G_{U}\tau^{P^{3}_{min}})$. It is clear that the algorithm is exponential in $\tau$ and also clearly worse (in terms of $\tau$) than the corresponding classical complexity given by $\mathcal{O}(\log(N_{g}s\tau\kappa\log(1/\varepsilon)))$ [harrow2009quantum, shewchuk1994introduction]. Again, although the algorithm is still logarithmic in $N_{g}$, it offers the poorest complexity scaling compared to all other TMCQCs presented here.
@@ IV.4 TMCQC4 - Implicit Serial Circuit
Gate complexity - The truncated series has a total of $\mathcal{O}(P^{3}_{min})$ terms, so the total depth can be estimated to be $\mathcal{O}(P^{3}_{min}\tau)$. Thus it can be seen readily that the total gate complexity is $\mathcal{O}\Big{(}s\epsilon,\log(N_{g}\tau),\textrm{polylog}(\epsilon/\varepsilon_{U},1/\varepsilon_{N}),\log^{-3}((\kappa-1)^{-1})\Big{)}$. Comparing this with the classical complexity [harrow2009quantum, shewchuk1994introduction] of $(\log(N_{g}s\tau\kappa\log(1/\varepsilon)))$, one finds that the algorithm (without any parallelization) has a comparable scaling in $\tau$, while it is exponentially better than classical in $N_{g}$ and $\kappa$.
@@ IV.6 Query complexity and success probability
The above value is for a single linear combination of unitaries block. However in essence, each TMCQC provides a different unitary block-encoding of the corresponding $A$ to perform time marching. The $p_{\textrm{succ}}$ of the solution at $t=\tau\Delta t$ thus varies for each TMCQC. The specific dependence arises from the difference in the number of times the unitaries oracle is queried and the corresponding depth, which we shall represent from now on as $G_{\textrm{L}}$. Comparing all the TMCQCs we can easily identify that in TMCQC2 and TMCQC4, the decay in $p_{\textrm{succ}}$ is particularly the most severe since the unitaries are applied in series. Therefore for $\tau$ such applications, the RHS of eq. (18) will now be raised to the power $\tau$. For generality, let us represent this exponent with $G_{\textrm{L}}$. A single application of the two-unitary oracle involves preparing a superposition state by Hadamard gates acting on ancilla qubits. This implies that $\beta=2$. Since the $p_{\textrm{succ}}$ for every time step is independent, the probability thus tends to decay exponentially [berry2024quantum, fang2023time], compounding for each of the $\tau$ steps. Apart from this, the overall success probability also depends on ratio of the norms of the initial and final time solution states. Thus, we can rewrite $p_{\textrm{succ}}$ as
@@
In summary, it is ideal to ensure that $\eta\lesssim 1$; otherwise it would lead to a large pre-factor $\mathcal{O}(2^{\log(\tau)/\log(\kappa-1)})$ at best, or $\mathcal{O}(2^{\tau^{2}})$ at worst. A note on condition number dependence is needed here. The discussion so far suggests only logarithmic dependence on the condition number when compared to earlier works that demonstrate an optimal result of linear $\mathcal{O}(\kappa)$ dependence [costa2022optimal]. Although this is indeed the case, it comes at a certain cost described below.
  • It is important to note the distinction between quantum linear solvers and time marching solvers (this work). Given this, unlike earlier work, we use a distinct LCU decomposition combined with a specific extrapolation technique that leads to a different complexity scaling. Also, as mentioned in Table IV.5, if an explicit amplitude amplification is required [childs2017quantum], we pick up a $\mathcal{O}(\kappa)$ factor. However, the question here is whether we could lower the required query complexity without an explicit amplitude amplification, though the use of an amplitude amplification step is not precluded.
  • The linear dependence on the condition number as shown in earlier optimal linear solvers is now transformed into a dependence on the quantities $\epsilon$ and $||\textbf{A}||$ in the following sense. The solution computed by the approximate inverse of the matrix system is dictated by the number of terms $P_{\text{min}}$, which in turn depends on $\kappa$ and $||\textbf{A}||$. We observe that the requirement of a bounded $\kappa$ does not stymie the range of system parameters that can be explored. But this transfers the onus onto the norm of a given matrix, $||\mathbf{A}||$. This quantity, however, can be scaled (by the factor $\delta$) and conditioned on this, the overall query complexity is subsequently scaled by a factor $\eta^{2G_{\textrm{L}}}=(1/\epsilon\delta||A||)^{2G_{\textrm{L}}}$ as mentioned in the main text in Section IV.F. This may be further improved by (say) quantum amplitude amplification to a factor $(1/\epsilon\delta||A||)^{G_{\textrm{L}}}$. For the case of TMCQC-5,6, where we apply the LCU decomposition once, we have $\eta=(1/\epsilon\delta||A||)$. This quantity determines the query complexity. This factor tends to be a large constant (corresponding to low probably solution subspace for post selection), thus implying a larger query complexity. Now, one may regard that the complexity of our algorithm scales linearly in $\eta$, which is analogous to the optimal result in terms of condition number $\kappa$, compared to earlier results.
  • We now mention an advantage here. Unlike the condition number that is controlled via relatively involved preconditioning techniques, the factor $\eta$ can now be lowered and successively made close to unity via iterative applications of Richardson’s extrapolation, since the parameter $\epsilon$ can be now controlled; this, in turn, avoids $\eta$ from being an unfavorably large constant.
  • In essence this work is an improvement (even if not exponential) over the previous optimal results in terms of condition number.
To see this, while ensuring that the query complexity is controllable, we need to pay special attention to the following three situations.
@@ V End-to-end algorithm and near-term strategies
(A) Shot parallelization - Generally the shot count available on real quantum hardware (here, IBMQ) is limited to about $\sim 3\times 10^{4}\textrm{~{}to~{}}5\times 10^{5}$ shots. If we consider a simulation requiring a large shot-count of about (say) $N_{s}=2^{20}$ shots or higher, such an execution would not be possible with current hardware. A simple workaround to achieve this would be actually to execute multiple quantum circuits in parallel. For the current example, naively, we can perform two parallel circuit executions with $2^{19}$ shots each. When possible, we could go even further by executing several low shot-count circuits in parallel, with which we can increase the total shot count while improving the sampling accuracy of the quantum state. It is important to note here that this implies a linear space-time trade-off, under the assumption of access to multiple qubits of similar quality (calibrated error rates) within the same device. This is not a fundamental assumption but the qubit quality varies in pratice across the device at any given instant of time. To fit our simulations using only the available shot count, we could instead use the Richardson extrapolation to lower the required $N_{s}$. This is discussed below in this section.
(B) LCU parallelization - Let us consider TMCQC5 and TMCQC6, the best circuits introduced here in terms of optimality. These systems encode a full history state, simulating the entire time evolution in one shot. The required number of ancilla qubits, thus the control qubit operations, depend logarithmically on the depth of LCU decomposition as outlined earlier in Section IV. LCU parallelization entails implementing these controlled unitaries in parallel, either as parallel and separate circuits or within the same circuit. For clarity, let us consider a single simulation time step, using either the two or four unitary block encodings.
@@
    Naive Parallel unitaries - We execute each unitary $U_{i}$ (without any controls) in parallel on 2 (or 4) separate circuits, starting with the same initial state. This requires $3n_{g}$ qubits in total. Since each controlled unitary is parallelized, the control qubits can now be discarded. This also implies that the resulting decomposition into two qubit gates will be significantly efficient in gate complexity. The number of parallel circuits, and thus the number gates, is determined by the LCU decomposition. As discussed in Section IV, this is logarithmic in system parameters and as a consequence results in gate complexity of the parallel circuits to be efficient. At most a single control qubit might be necessary based on how the partial solutions are combined subsequently. When the resulting parallel (partial) solution states have purely real amplitudes, we can perform a straightforward measurement in the computational basis of each partial solution. However, the sign of negative numbers requires additional measurements, although not a full tomography. Following this, these partial solutions can be combined classically to output the final solution. However, if any of the partial solutions have complex amplitudes, a direct extraction of solutions requires expensive tomography to reconstruct the state. Two possibilities exist. In the first case, we combine those parallel unitaries as a single circuit such that the grouping produces real amplitudes for the partial states. From here, we can proceed with measurement as earlier. This new grouping would, of course, increase the depth of the circuit, which is now at some intermediate level of parallelization. Secondly, if we are interested only in computing the expectation value of a certain observable, instead of the full state, it can be done while still retaining the full parallelization. The expectation value of an operator $U_{i}$ can be computed accurately by using a Hadamard Test circuit. A direct Hadamard test outputs $Re(\langle\psi|U_{i}|\psi\rangle)$. For the imaginary part $Im(\langle\psi|U_{i}|\psi\rangle)$, a simple modification is done by adding the $S^{{\dagger}}$ gate as shown in figure 2(b).
@@
    Fanout Parallel unitaries - A more robust parallelization is possible by invoking the idea of fanout quantum circuits [nielsen2010quantum, cleve2000fast, moore2001parallel], though it is hard to realize using near-term devices. Here, the circuit can be parallelized by a single entangled quantum circuit (not separate circuits as earlier) at the cost of extra ancilla qubits [hoyer2005quantum, boyd2023low, zhang2024parallel]. In these types of strategies, the state of a single ancilla register prepared with the target state is then “basis-type copied” (fanout) to other additional ancilla registers by applying controlled CNOT gates as $|\psi\rangle|0\rangle\cdots|0\rangle\rightarrow|\psi\rangle|\psi\rangle\cdots|\psi\rangle$. Now the unitary operators are applied in parallel on copies of the target state $|\psi\rangle$. Especially when the Hamiltonian that is being simulated has certain properties [zhang2024parallel], or when the set of unitaries can be partitioned into Pauli operators, such fanout schemes can be used, with the aid of Clifford circuits [boyd2023low], to parallelize the overall algorithm.
@@
where $\epsilon_{1}>\epsilon_{2}$ and $\gamma=\epsilon_{1}/\epsilon_{2}$ is the order of extrapolation, where the error of extrapolation scales as $\mathcal{O}(\epsilon^{4})$. $\Gamma$ here represents, collectively, any arbitrary set of parameters on which the operator could depend. Higher powers of $\gamma$ lead to higher orders and more accurate extrapolations. As we shall demonstrate later through simulations, even for $\epsilon_{1}and\epsilon_{2}$ close to unity, the solution can be computed accurately through extrapolation. This method could also serve as an aid to possible amplitude amplification procedures [berry2014exponential]. The extrapolation procedure can be viewed alternatively as follows. Given a fixed number of shots $N_{s}$, we can extrapolate the solutions that were obtained with $\epsilon_{1,2}$ and $N_{s}$ shots as